OpenStudy (anonymous):

SENIOR IN HIGH SCHOOL NEED HELP ASAP Jenny is plans to have two children but doesn't know if they are boy-boy, girl-girl, or boy-girl. What is the probability that she will have two boys?

OpenStudy (anonymous):

okay, your sample space (all possible outcomes) for this are the combinations of boy and girl, (boy,girl) or (girl,boy) or (girl,girl) or (boy,boy) now remember that probability is the number of favorable outcomes, meaning how many times what you want occurs, divided by the total amount of outcomes that could happen, so, now i'll ask you, what's the probability that you get the result (boy,boy)

OpenStudy (anonymous):

25 percent

OpenStudy (tkhunny):

Horrible question. It's a "plan", not a reality. All this must factor in the probability of No Children, 1 Child, and 2 Children. Apparently, she intends to give away any others after 2 . Seriously, this is a horrible question.

OpenStudy (anonymous):

Haha, when you point that out, yeah it sounds horrible, but then again, when i personally read these questions, i expect them to be straight forward, they are in the Mathematics section after all, not the Literature

OpenStudy (tkhunny):

Tell that to all the folks who would like to have children and can't manage. They already feel marginalized in some portions of our society. Anyway, GIVEN that there are two children, AND that P(Boy) = P(Girl) = 0.50, we have it as you stated. p(Boy, Boy) = 0.5*0.5 = 0.25 p(Girl, Girl) = 0.5*0.5 = 0.25 p(Boy and Girl) = p(Boy,Girl)+p(Girl,Boy) = 0.5*0.5 + 0.5*0.5 = 0.25+0.25 = 0.50